# TabKAN: Advancing Tabular Data Analysis using Kolmograv-Arnold Network

### v1

@@ 3.9 Rational Kolmogorov-Arnold Network (RKAN)
$\mathbf{y}=\sum_{n=0}^{N}\theta_{n}\,J_{n}^{(\alpha,\beta)}(\phi(\mathbf{x},L))$
@@
RKAN leverages Optuna to adaptively choose the network depth, layer widths, and polynomial orders, and employs the L-BFGS optimizer to minimize cross-entropy loss. By combining rational mapping with Jacobi polynomials, RKAN effectively models asymptotic behavior and achieves robust performance across various datasets, as evidenced by metrics such as F1 score and ROC-AUC.
@@ 4 Experiments and Results
We thoroughly evaluate our model using ten publicly available datasets covering supervised and transfer learning scenarios. Although we use several performance metrics—AUC, F1 score, precision, and recall—we report only AUC due to its effectiveness in summarizing classification performance and space constraints. To assess robustness, we benchmark our model against state-of-the-art methods with varying data and feature usage, using average ranking as the primary comparison metric, following the protocol of [13], which offers a holistic view of relative performance across datasets.
@@ 4.1 Datasets
Collectively, these benchmark datasets span diverse fields and data structures, enabling a thorough assessment of our approach. Additional details for each dataset appear in Table 1.
@@
We choose the configuration that yields the highest validation performance and then train the model on each dataset using ten distinct random seeds to mitigate the impact of training variability. This procedure aligns with the comparative approach used in TabMixer [24]. All experiments run on an AMD Ryzen Threadripper PRO 5965WX 24-core CPU with 62 GB of RAM and an NVIDIA RTX A4500 GPU featuring 20 GB of memory.
@@ 4.3 Supervised Learning
Table 2 presents results from evaluating classical baselines (Logistic Regression, XGBoost, MLP, SNN), specialized tabular models (TabNet, DCN, AutoInt, TabTrans, FT-Trans, VIME, SCARF, TransTab), and a range of KAN variants (ChebyshevKAN, JacobiKAN, PadeRKAN, Fourier KAN, fKAN, fast-KAN) alongside the original KAN architecture. Each approach is optimized individually, with hyperparameters tuned to account for differences in model structure and dataset characteristics. The results shows Chebyshev KAN emerges as the top performer, suggesting that its Chebyshev polynomial basis effectively captures intricate decision boundaries while maintaining stability. However, certain KAN variants, such as wav-KAN [36], did not show strong performance on any dataset, while others, like fc-KAN [37], feature a heavier architecture that limits practical applicability. This highlights the inherent difficulty of designing a single architecture that performs uniformly well across diverse data distributions, leading us to exclude them from our final results.
@@ 4.4 Transfer Learning
Transfer learning enables machine learning models to leverage knowledge acquired from a source task to improve performance on a related target task through fine-tuning. While this paradigm has proven effective in domains like computer vision and natural language processing, where shared structures simplify knowledge transfer, its application to tabular data presents unique challenges. Feature heterogeneity, dataset-specific distributions, and the absence of universal structural patterns often lead to encoder overspecialization during conventional supervised pretraining. In this scenario, models trained on classification objective develop representations overly tailored to dominant patterns in the source data, which can limit their adaptability to target tasks with divergent feature spaces, class imbalances, or objectives.
@@
To systematically evaluate transfer learning in tabular settings, we adopt the methodology of [13], partitioning each dataset into two subsets (Set1 and Set2) with a controlled 50% feature overlap. This design introduces a domain shift by treating overlapping features as shared knowledge and non-overlapping features as distinct statistical domains. The partial overlap allows assessment of the model’s ability to generalize shared representations while adapting to novel features.
@@
This process involves two primary stages: pretraining and fine-tuning. During pretraining, the model is trained on Set1 using supervised learning to acquire initial feature representations. Upon convergence, the learned patterns are preserved by freezing all layers except the final main layer (and bias layer, if applicable). Subsequently, the model undergoes fine-tuning on Set2, where only the unfrozen layers are updated. This selective updating allows the model to adapt to the target data distribution of Set2 while retaining the knowledge acquired from Set1.
@@
To validate robustness, we evaluate the model on the test split of Set2, measuring its capacity to generalize under domain shifts. Additionally, we reverse the roles of Set1 and Set2 in a cross-validation framework, ensuring comprehensive evaluation of bidirectional knowledge transfer. This approach enables TabKAN to balance representation stability (via frozen layers) and adaptability (via targeted fine-tuning), addressing the challenges of feature heterogeneity and encoder overspecialization in tabular data.
The results shows that Fourier KAN outperforms all other architectures, suggesting that its kernel approximation via Fourier series provides a strong inductive bias for knowledge transfer. The smooth, periodic basis functions in Fourier expansions help capture recurring patterns in the data, thereby enabling more effective generalization to new domains.
@@ 4.5 Multi-class Classification
Table 4 compares TabKAN with other neural network models on two multi-class classification datasets. Because these tasks often suffer from sample imbalance, we used macro-F1 as the primary evaluation metric during training to ensure balanced performance across all classes [41]. All KAN variants consistently outperform the baselines, with JacobiKAN achieving the best overall results. Its use of Jacobi polynomials, which include additional parameters $\alpha$ and $\beta$ for shaping the polynomial basis, enables more flexible and precise kernel approximations, thereby capturing complex patterns more effectively.
@@ 4.6 Interpretability
For interpretability purposes, there are two main approaches: model-specific methods and model-agnostic methods. Model-specific methods apply only to a particular model (e.g., in logistic regression, coefficients represent feature importance). In contrast, model-agnostic methods can be applied to any model, regardless of its type.
A defining strength of Kolmogorov-Arnold Networks (KAN) lies in their inherent interpretability, a critical advantage for domains requiring transparency in decision-making. Unlike traditional black-box models (e.g., deep neural networks or gradient-boosted trees), KANs explicitly parameterize univariate functions along network edges, enabling direct visualization of feature-wise contributions and functional mappings. This architectural design not only reduces reliance on post hoc interpretation methods but also aligns model behavior with domain-specific knowledge, fostering trust and facilitating model audits.
Figures 1, 2, and 3 illustrate the attribution of Feature 1 in the CA dataset towards the output, while Figures 4, 5, and 6 represent the attribution of Feature 2. Specifically, Figures 1 and 4 demonstrate the interpretability of the Fourier-based KAN, whereas Figures 2 and 5 highlight the interpretability of the Chebyshev-based KAN. The domain and codomain are normalized for these figures, which explains the difference in scale compared to the PDP figures. We employed Partial Dependence Plot (PDP) as a post-hoc analysis to assess the differences in feature interpretation.
In our framework, each feature’s learned univariate function provides non-parametric insights into its contribution to the final prediction. By visualizing these functions, we observe how individual features nonlinearly transform inputs, revealing monotonic trends, thresholds, or saturation effects that may correlate with domain expertise.
The model’s structure also facilitates analysis of feature interactions. While KANs primarily model univariate functions, multivariate interactions emerge implicitly through the additive composition of these functions in deeper layers. By analyzing co-variations in learned functions across related features, we infer latent dependencies and validate them against domain knowledge. For example, overlapping saturation points in two features’ functions may indicate synergistic effects.
Figure 1: Fourier KAN - Feature 1
Figure 2: ChebyshevKAN - Feature 1
@@
Figure 4: Fourier KAN - Feature 2
Figure 5: ChebyshevKAN - Feature 2
@@ 4.7 Feature Importance and Dimensionality Reduction
Figure 7: Feature Importance
Figure 8: Feature Importance
@@
Figure 9: AUC vs Percentage of Top Selected Features
Figure 10: AUC vs Percentage of Top Selected Features
@@ 5 Ablation Study
5.1 Fine-tunning:
In transfer learning scenarios, where a pre-trained model is adapted to a new task or domain, the Group Relative Policy Optimization (GRPO) [42] framework provides a robust mechanism for fine-tuning by strategically balancing task-specific adaptation with knowledge retention. Leveraging a policy gradient approach, GRPO optimizes model parameters $\theta$ through advantage-weighted updates derived from reward signals ($R\in\{0,1\}$), where rewards reflect alignment between sampled predictions ($o\sim\pi_{\theta}$) and ground-truth labels. To mitigate catastrophic forgetting—a common challenge in transfer learning—the method incorporates a Kullback-Leibler (KL) divergence penalty $\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$, which constrains deviations from the reference policy $\pi_{\text{ref}}$ (e.g., the original pre-trained model). By sampling $G$ candidate predictions per input and computing normalized advantages $\hat{A}=R-\mathbb{E}[R]$, GRPO promotes exploration while maintaining stability, making it particularly suited for scenarios with limited target-domain data.
@@ 6 Conclusion
In this work, we introduced TabKAN, a novel Kolmogorov-Arnold Network-based architecture designed to advance tabular data analysis. By leveraging modular KAN-based models, we demonstrated their effectiveness in both supervised learning and transfer learning, where TabKAN significantly outperformed classical and Transformer-based models in knowledge transfer tasks. Unlike conventional deep learning approaches that rely on post hoc interpretability methods, TabKAN introduces model-specific interpretability, allowing for direct visualization and analysis of feature interactions within the network. To further enhance the expressiveness and adaptability of KANs for tabular data learning, we utilized multiple specialized architectures, including ChebyshevKAN, JacobiKAN, PadeRKAN, FourierKAN, fKAN, and fast-KAN, each offering distinct advantages in function approximation and performance. Additionally, we introduced a new fine-tuning approach for transfer learning, improving the efficiency of knowledge transfer between domains. Furthermore, we enhanced the basic architecture by leveraging the advantages of KAN over MLP, leading to the development of more robust and scalable models for tabular data. Our extensive experiments across multiple benchmark datasets highlight the robustness, efficiency, and scalability of KAN-based architectures, bridging the gap between traditional machine learning and deep learning for structured data. Building on these advancements, future work will focus on further optimizing KAN architectures and extending their applicability to self-supervised learning and domain adaptation, enabling broader adoption of KANs in real-world applications.
@@ Appendix A Hyperparameter Sensitivity
This appendix provides a detailed analysis of the hyperparameter sensitivity across seven neural network models (fastKAN, rKAN, fKAN, ChebKAN, KAN, FourierKAN) evaluated on eight datasets (IO, IC, DS, CG, CB, CA, BL, AD). The analysis focuses on four key architectural metrics: Layers, Neurons, Order, and Grid, as summarized in Tables LABEL:tbl:h_io,_tbl:h_ic,tbl:h_ds,tbl:h_cg,tbl:h_cb,tbl:h_ca,tbl:h_bl,tbl:h_ad.
@@
Appendix B Consistency Across 100 Seed Values
Our experiments reveal that the choice of seed plays a crucial role in influencing the results. This effect is particularly evident during the data partitioning process for training and testing. To capture this variability, we illustrate the interquartile range, offering a broader perspective on the fluctuations in our findings, as depicted in Fig. 12. This analysis highlights both the stability of our model and the inevitable variations in performance stemming from different data splits.
Figure 12: The interquartile range across 100 runs with varying seed values highlights the influence of random seed selection on experimental outcomes. The plot depicts the variation in the distribution of raw and synthesized data across different training and test set splits.

### v2

@@ 3.2.7 Rational Kolmogorov-Arnold Network (RKAN)
$\mathbf{y}=\sum_{n=0}^{N}\theta_{n}\,J_{n}^{(\alpha,\beta)}(\phi(\mathbf{x},L)),$
@@
3.3 Neural Architecture Search
Neural Architecture Search (NAS) is the automated process of identifying optimal neural network configurations for a specific learning task. It replaces manual architecture design with a systematic exploration of hyperparameters to improve model performance. In our study, we employ the Optuna framework [31], which efficiently navigates the search space through Bayesian optimization and pruning strategies. For each KAN variant, we perform a dedicated NAS procedure to determine the most effective combination of architecture and functional parameters.
For FastKAN, we tune the number of layers $L$, the width vector $\mathbf{w}=(w_{1},\ldots,w_{L})$, and the parameters of the RBF activation functions. In PadéRKAN, we optimize network depth, layer widths, and the polynomial degrees $(q,k)$. For FourierKAN, the grid size $g$, which controls the frequency resolution of the Fourier expansion, is selected through NAS. The fKAN model includes hyperparameters such as depth, widths, and the Jacobi polynomial order. Finally, RKAN uses NAS to select depth, widths, and Jacobi polynomial order to adapt the rational architecture to varying dataset complexities.
All models are trained using the L-BFGS optimizer with cross-entropy loss. Validation F1 score is used as the selection criterion, ensuring that the chosen configurations generalize well while adapting to specific structural and statistical properties of the data. The results and analysis of these hyperparameter optimization procedures are reported in Appendix A.
@@ 4 Experiments and Results
We evaluate our model on ten publicly available datasets across both supervised and transfer learning tasks. While multiple performance metrics are computed—AUC, F1 score, precision, and recall—we report only AUC due to its effectiveness in summarizing classification performance and limitations on space. To assess robustness, we compare our model with state-of-the-art baselines under varying data and feature configurations. Following the protocol in [14], we use average ranking as the main comparison criterion, which provides an overall view of relative performance across datasets.
@@ 4.1 Datasets
We choose the configuration that yields the highest validation performance and then train the model on each dataset using ten distinct random seeds to mitigate the impact of training variability. This procedure aligns with the comparative approach used in TabMixer [25]. All experiments run on an AMD Ryzen Threadripper PRO 5965WX 24-core CPU with 62 GB of RAM and an NVIDIA RTX A4500 GPU featuring 20 GB of memory. To improve inference efficiency while preserving accuracy, we used PyTorch’s torch.quantization package to implement both static and dynamic post-training quantization, as well as quantization-aware training (QAT) [32]. This reduced the memory footprint of some models by 3% to 15%, without a significant loss in accuracy.
@@ 4.3 Supervised Learning
Table 2 presents results from evaluating classical baselines (LR, XGBoost, MLP, SNN), specialized tabular models (TabNet, DCN, AutoInt, TabTrans, FT-Trans, VIME, SCARF, TransTab), and a range of KAN variants (ChebyshevKAN, JacobiKAN, PadeRKAN, Fourier KAN, fKAN, fast-KAN) alongside the original KAN architecture. Each approach is optimized individually, with hyperparameters tuned to account for differences in model structure and dataset characteristics. The results shows Chebyshev KAN emerges as the top performer, suggesting that its Chebyshev polynomial basis effectively captures intricate decision boundaries while maintaining stability. However, certain KAN variants, such as wav-KAN [38], did not show strong performance on any dataset, while others, like fc-KAN [39], feature a heavier architecture that limits practical applicability. This highlights the inherent difficulty of designing a single architecture that performs uniformly well across diverse data distributions, leading us to exclude them from our final results.
@@ 4.4 Transfer Learning
Transfer learning permits machine learning models to leverage knowledge acquired from a source task to enhance performance on a related target task through fine-tuning. Although this paradigm has proven effective in domains such as computer vision and natural language processing, where shared structures simplify knowledge transfer, its application to tabular data presents unique challenges. Feature heterogeneity, dataset-specific distributions, and the absence of universal structural patterns often lead to encoder overspecialization in conventional supervised pretraining. In this scenario, models trained with a classification objective develop representations that become overly specific to dominant patterns in the source data, which restricts their adaptability to target tasks with divergent feature spaces, class imbalances, or differing objectives.
@@
To systematically evaluate transfer learning in tabular settings, we adopt the methodology of [14] and partition each dataset into two subsets, Set1 and Set2, with a controlled 50% feature overlap. This design introduces a domain shift by treating overlapping features as shared knowledge while regarding non-overlapping features as distinct statistical domains. The partial overlap allows an assessment of the model’s ability to generalize shared representations while adjusting to novel features.
@@
The process involves two primary stages: pretraining and fine-tuning. During pretraining, the model receives supervised training on Set1 to acquire initial feature representations. Once convergence is achieved, all layers except for the final main layer (and the bias layer, if applicable) are frozen, thereby preserving the learned patterns. Subsequently, the model undergoes fine-tuning on Set2, with only the unfrozen layers being updated. This selective updating permits the model to adapt to the target data distribution of Set2 while retaining the knowledge acquired from Set1.
@@
Robustness is validated by evaluating the model on the test split of Set2 to measure its capacity to generalize under domain shifts. In addition, we reverse the roles of Set1 and Set2 in a cross-validation framework, which ensures a comprehensive assessment of bidirectional knowledge transfer. This approach allows TabKAN to balance stability in feature representations with adaptability through targeted fine-tuning, thereby addressing the challenges posed by feature heterogeneity and encoder overspecialization in tabular data.
As shown in Table 3, Fourier KAN attains the highest average performance among all KAN-based architectures (0.859) and ranks second overall across all evaluated models, outperforming both Transformer-based and classical approaches. Its Fourier series expansion proves particularly effective for transfer learning. The smooth, periodic basis functions approximate both low- and high-frequency components, which supports adaptation to varying feature distributions across source and target domains.
Other KAN variants such as JacobiKAN (0.814), ChebyshevKAN (0.796), and the base KAN model (0.774) also achieve strong results, often outperforming standard baselines such as XGBoost (0.776), MLP (0.775), and Transformer-based methods including TabTransformer (0.764), AutoInt (0.754), and DCN (0.758).
@@ 4.5 Multi-class Classification
Table 4 presents a comparison between TabKAN and several neural network baselines on two multi-class classification benchmarks. Since these tasks often involve class imbalance, macro-F1 was selected as the primary evaluation metric during training to ensure balanced performance across all classes [43]. All KAN variants consistently outperform baseline models, with JacobiKAN achieving the highest overall performance. Its use of Jacobi polynomials, parameterized by $\alpha$ and $\beta$, provides a more adaptable polynomial basis, which supports improved approximation of complex patterns.
@@ 4.6 Interpretability
For interpretability purposes, there are two main approaches: model-specific methods and model-agnostic methods. Model-specific methods apply only to a particular model (e.g., in LR, coefficients represent feature importance). In contrast, model-agnostic methods can be applied to any model, regardless of its type.
A defining strength of KAN lies in their inherent interpretability, a critical advantage for domains requiring transparency in decision-making. Unlike traditional black-box models (e.g., deep neural networks or gradient-boosted trees), KANs explicitly parameterize univariate functions along network edges, enabling direct visualization of feature-wise contributions and functional mappings. This architectural design not only reduces reliance on post hoc interpretation methods but also aligns model behavior with domain-specific knowledge, fostering trust and facilitating model audits.
Figures 1, 2, and 3 illustrate the attribution of Feature 1 in the CA dataset towards the output, while Figures 4, 5, and 6 represent the attribution of Feature 2. Specifically, Figures 1 and 4 demonstrate the interpretability of the Fourier-based KAN, whereas Figures 2 and 5 highlight the interpretability of the Chebyshev-based KAN. The domain and co-domain are normalized for these figures, which explains the difference in scale compared to the Partial Dependence Plot (PDP) figures. We employed PDP as a post-hoc analysis to assess the differences in feature interpretation.
In our framework, each feature’s learned univariate function provides non-parametric insights into its contribution to the final prediction. Visualization of these functions reveals how individual features transform inputs in nonlinear ways, uncovering monotonic trends, thresholds, or saturation effects that may align with known domain behavior.
The model’s structure also supports investigation of feature interactions. Although KANs model univariate functions explicitly, deeper layers combine these mappings additively, allowing complex multivariate dependencies to emerge. Co-variations across learned functions of related features can indicate latent interactions, which are then compared with domain knowledge. For instance, overlapping saturation patterns in multiple feature mappings may reflect synergistic influence.
Figure 1: Attribution of Feature 1 toward the output prediction using the Fourier KAN
Figure 2: Attribution of Feature 1 toward the output prediction using the ChebyshevKAN
@@
Figure 4: Attribution of Feature 2 toward the output prediction using the Fourier KAN
Figure 5: Attribution of Feature 2 toward the output prediction using the ChebyshevKAN
@@ 4.7 Feature Importance and Dimensionality Reduction
Figure 7: Feature Importance based on ChebyshevKAN
Figure 8: Feature Importance based on Fourier KAN
@@
Figure 9: AUC vs Percentage of Top Selected Features for Fourier KAN
Figure 10: AUC vs Percentage of Top Selected Features for ChebyshevKAN
@@ 5 Ablation Study
5.1 Fine-tunning
In transfer learning scenarios, where a pre-trained model is adapted to a new task or domain, the Group Relative Policy Optimization (GRPO) [44] framework provides a robust mechanism for fine-tuning by balancing task-specific adaptation and knowledge retention. Using a policy gradient method, GRPO optimizes model parameters $\theta$ through advantage-weighted updates derived from reward signals ($R\in\{0,1\}$), which measure the alignment between sampled predictions ($o\sim\pi_{\theta}$) and ground-truth labels. To address catastrophic forgetting-a typical issue in transfer learning-the method includes a Kullback-Leibler (KL) divergence penalty $\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$, which constrains deviations from the reference policy $\pi_{\text{ref}}$ (e.g., the original pre-trained model). By sampling $G$ candidate predictions per input and calculating normalized advantages $\hat{A}=R-\mathbb{E}[R]$, GRPO promotes exploration while maintaining stability, making it well-suited for tasks with limited target-domain data.
@@ 6 Conclusion
In this work, we introduced TabKAN, a novel Kolmogorov-Arnold Network-based architecture designed to advance tabular data analysis. By leveraging modular KAN-based models, we demonstrated their effectiveness in both supervised learning and transfer learning, where TabKAN significantly outperformed classical and Transformer-based models in knowledge transfer tasks. Unlike conventional deep learning approaches that rely on post hoc interpretability methods, TabKAN introduces model-specific interpretability, allowing for direct visualization and analysis of feature interactions within the network. To further enhance the expressiveness and adaptability of KANs for tabular data learning, we utilized multiple specialized architectures, including ChebyshevKAN, JacobiKAN, PadeRKAN, FourierKAN, fKAN, and fast-KAN, each offering distinct advantages in function approximation and performance. Additionally, we introduced a new fine-tuning approach for transfer learning, improving the efficiency of knowledge transfer between domains. Furthermore, we enhanced the basic architecture by leveraging the advantages of KAN over MLP, leading to the development of more robust and scalable models for tabular data. Our experiments across multiple benchmark datasets highlight the robustness, efficiency, and scalability of KAN-based architectures, bridging the gap between traditional machine learning and deep learning for structured data. Building on these advancements, future work will focus on further optimizing KAN architectures and extending their applicability to self-supervised learning and domain adaptation, enabling broader adoption of KANs in real-world applications.
7 Acknowledgements
We thank the creators of the public datasets and the authors of the baseline models for making these resources available for research. We gratefully acknowledge Brian Gold, PhD, and the Gold Lab at the University of Kentucky for their support and for providing the facilities necessary to carry out this research. This research is supported in part by the NSF under Grant IIS 2327113 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01.
@@ Appendix A Hyperparameter Sensitivity
This appendix provides a detailed analysis of the hyperparameter sensitivity across seven neural network models (fastKAN, RKAN, fKAN, ChebKAN, KAN, FourierKAN) evaluated on eight datasets (IO, IC, DS, CG, CB, CA, BL, AD). The analysis focuses on four key architectural metrics: Layers, Neurons, Order, and Grid, as summarized in Tables 6, 7, 8, 9, 10, 11, 12, and 13.
@@
Figure 12: The interquartile range across 100 runs with varying hyperparameters highlights the influence of architecture on experimental outcomes. The plot depicts the variation in the distribution of raw and synthesized data across different training and test set splits.
Appendix B Dataset links
We provide the links for the public datasets that we used for the benchmark. Details of each dataset can be found in Table 14.
Table 14: Benchmark Dataset Links
Dataset URL
Credit-G https://www.openml.org/search?type=data&status=active&id=31
Credit-Approval https://archive.ics.uci.edu/ml/datasets/credit+approval
Dress-Sales https://www.openml.org/search?type=data&status=active&id=23381
Adult https://www.openml.org/search?type=data&status=active&id=1590
Cylinder-Bands https://www.openml.org/search?type=data&status=active&id=6332
Blastchar https://www.kaggle.com/datasets/blastchar/telco-customer-churn
Insurance-Co https://archive.ics.uci.edu/ml/datasets/Insurance+Company+Benchmark+(COIL+2000)
1995-Income https://www.kaggle.com/datasets/lodetomasi1995/income-classification
ImageSegmentation https://www.openml.org/search?type=data&sort=version&status=any&order=asc&exact_name=segment&id=36
ForestCovertype https://archive.ics.uci.edu/dataset/31/covertype